OpenStudy (anonymous):

area of the surface defined by x+y+z=1, x^2+2y^2<1

OpenStudy (anonymous):

using calculus?

OpenStudy (anonymous):

using green or stokes theorem in multivariable calculus

OpenStudy (dumbcow):

the way i see it is this is area of surface of inside of ellipse which lies on a plane equation of plane: z = 1-x-y limits of x,y due to ellipse -1<x<1 |y|< sqrt((1-x^2)/2) \[\rightarrow \int\limits_{-1}^{1}\int\limits_{-\sqrt{(1-x^{2})/2}}^{\sqrt{(1-x^{2})/2}} (1-x-y) \sqrt{\frac{dz}{dx}^{2} +\frac{dz}{dy}^{2}+1} dy dx\] dz/dx = -1 dz/dy = -1

OpenStudy (dumbcow):

i am not familiar with greens or stokes thm though so i could be doing this wrong